Problem 10-7 Calculating Salvage Value [LO1] Consider an asset that costs $720,000 and is depreciated straight-line to zero over its eight-year tax life. The asset is to be used in a five-year project; at the end of the project, the asset can be sold for $159,000. If the relevant tax rate is 22 percent, what is the aftertax cash flow from the sale of this asset? (Do not round intermediate calculations.) Aftertax salvage value
Answer #1

Annual depreciation=(Cost-Salvage value)/Useful Life

=(720,000/8)=$90,000/year

Hence book value as on date of sale=Cost-Accumulated depreciation

=720,000-(90000*5)=$270,000

Hence loss on sale=(270,000-159000)=$111,000

Hence aftertax salvage value=Sale proceeds+(loss on sale*Tax rate)

=159000+(0.22*111000)

=$183,420.